OpenStudy (anonymous):

What are the points of intersection of y=abs(x^2-4) and y=(x^2/2)+4

OpenStudy (anonymous):

I hope you feel the shame that I feel radiating from this question. It pains me to ask an algebra question T_T

OpenStudy (mathstudent55):

\(|x^2 - 4| = \dfrac{x^2}{2} + 4\) \(x^2 - 4 = \dfrac{x^2}{2} + 4\) or \(-(x^2 - 4) = \dfrac{x^2}{2} + 4\) Solve for x. Then plug in the x values to find the y values.

OpenStudy (anonymous):

Oh right. Thanks!

OpenStudy (anonymous):

Refer to the attached Mathematica 9 plot.

OpenStudy (triciaal):